# Position: We need responsible, application-driven (RAD) AI research

### v1

@@ 1 Introduction
The artificial intelligence (AI) landscape is evolving rapidly, requiring researchers to engage with many different aspects of AI integration into society. Innovations span diverse fields, from specific examples in medicine - where AI is helping with early detection of atrial fibrilliation and providing real-time health insights in clinical settings (Briganti & Le Moine, 2020), to agriculture - where it promises to better predict crop yield (Elbasi et al., 2022), optimise fertiliser use (Elbasi et al., 2022), and analyse long-term cropland changes (Hartman et al., 2022). Across both the public and private sectors, the widespread uptake of AI elevates the need for applied AI research to be responsive to contextual needs and societal values (Nissenbaum, 2009), ethical and legal considerations, and public concerns. Further, AI researchers have responsibilities to inform and educate the public regarding actual AI capabilities and limitations, as an instrumental component of building trust in human-AI collaboration (McGrath et al., 2024).
@@
In recent years, the concept of Responsible AI (RAI) has gained significant traction in industry, government, and academic settings (Adams et al., 2024; Maslej et al., 2024; UNESCO, 2024). While RAI is not a novel concept (Confalonieri et al., 2021; Lu et al., 2023), it has become omnipresent as AI capabilities have become embedded into everyday technologies and the failures, flaws, and harms of AI have been made visible (Cressie, 2023). Responsible AI draws on prior technical work undertaken in the areas of fairness, transparency, and explainability, as means of responding to AI’s deficiencies (Confalonieri et al., 2021). Yet, two of the major weaknesses of RAI are that it lacks stable definition and that it fails, at core, to specify who is responsible, for what, to whom (Powles, 2023). Instead, RAI casts responsibility diffusely. Adams et al. (2024), for example, define RAI as “the design, development, deployment and governance of AI in a way that respects and protects all human rights and upholds the principles of AI ethics through every stage of the AI lifecycle and value chain. It requires all actors involved in the national AI ecosystem to take responsibility for the human, social and environmental impacts of their decisions”.
@@
RAI gestures to earlier frameworks designed to ensure science and innovation is responsive to societal needs, including work on ethical, legal, and social implications, value sensitive design and, most significantly, Responsible Innovation (RI) or Responsible Research and Innovation (RRI). The classic formulation of RI advocates for a deliberative approach to research, centred on four dimensions: anticipation; reflexivity; inclusion; and responsiveness (Owen et al., 2013; Stilgoe et al., 2020). This model has been extensively applied and refined since the early 2010s, and adds considerable sophistication to RAI discourse, particularly around integrating affected communities into the research process (McCrea et al., 2024).
@@
This paper advocates for people-centered, participatory approaches in RAD-AI research, fostering collaboration, co-design, and efficacy-driven innovation beyond the constraints of purely methods-driven approaches. By embedding AI development within real-world contexts and diverse perspectives, we can ensure that these technologies are not only technically robust but also socially and ethically viable. The question is: how do we get there?
@@ 2 From AI to RAD-AI: Framing AI Research as a Sociotechnical Challenge
2.1 Paradigms: From Technological Determinism to Centering People
The AI community needs to be wary of adopting the paradigm of technological determinism: the theory that technology inherently and autonomously changes society, unaffected by social pressures and constructs (Kline, 2015). Technological determinism can lead to AI researchers assuming that their work is inherently desirable, with the affordances of technology “becom[ing] ends in themselves, instead of the means for creating desired social ends” (Kline, 2015). Instead of assuming, researchers should ask critical questions to ensure that their work does indeed align with contextual needs and societal values, shaping how technology functions, for whom, and to what effect (Davis, 2020). As Stilgoe and Cohen (2021) argue, by its nature, technological determinism precludes co-creation between those developing technologies and affected communities, since communities are the assumed eventual adopters of technology, rather than active players in how it is researched, developed, and shaped (Stilgoe & Cohen, 2021). The paradigm sees communities as “onlookers… displace[d] from the action” (Zenz & Powles, 2024), incorrectly assuming that adoption is inevitable and public approval is not required (Stilgoe & Cohen, 2021; Zenz & Powles, 2024).
@@ 2.2 Breaking Out of AI (Black) Boxes
Problem Framing. One challenge we identify is the boundaries that experts impose when defining problems. Traditionally, professionals such as doctors, engineers, lawyers, and computer scientists, are taught to “bound, map, characterize, and document” a problem, situating it in a specific technical framing (Li, 2007). To distill problems into technical solutions, the socio-political complexities that define them are mostly glossed over, creating a skewed and incomplete approach to knowledge and in some cases even a “sidelining of the data so painstakingly collected” (Li, 2007). While expertise is important, it may narrow the frame of reference and leave out critical components of the complex world. What scientists initially hypothesise, collect data on, and build models around reflects existing power dynamics and interests, worldviews, and technical training, potentially missing critical nuances. We may miss understanding and nuance because it is not captured in the original framing of the question, the data collected, or the type of result being prioritised, including the specific metrics used to measure success (Rolnick et al., 2024). Addressing real-world problems with algorithms and AI systems requires looking beyond the constraints experts impose - both in how challenges are framed and in the broader contexts that shape them. These contexts, spanning history, politics, and power (Powles, 2023), not only define but also limit potential solutions.
@@
A second challenge is where the expert box is also bound by scientific reductionism – the scientist’s proclivity for oversimplification of reality (Greenhalgh, 2025). Concerns about oversimplification of AI approaches are evident across various fields, ranging from medicine (Greenhalgh, 2025; Sturmberg & Mercuri, 2024) to agriculture (Tzachor et al., 2022). For example, model emulation approaches (MacKinlay et al., 2025; Bolt et al., 2023; Gladish et al., 2018), such as those applied to complex agricultural models like APSIM (Powell et al., 2023) offer crucial speedups and simplifications, effectively ‘jailbreaking’ their complexity. However, skepticism toward AI innovations often stems from the notion that AI seeks to replace, rather than complement and build off, well-established agricultural models. This kind of reductionism - which can involve excessive as well as insufficient parameterisation - generally overlooks the nuanced, interconnected, interdependent, and evolving nature of problems. It can also hide assumptions. As Sturmberg and Mercuri (2024) note, “no problem exists in isolation” rather “all problems are embedded within a unique context.” Similarly, RAD-AI approaches will have nuanced differences depending on the context in which they are applied.
@@
Method Selection. A key challenge in AI method selection is the lack of interoperability of complex models. While ML approaches such as neural networks, support vector machines, random forests, and k-nearest neighbours can be tuned for high accuracy, they often obscure intermediate decisions, making it difficult to understand how conclusions are reached (Confalonieri et al., 2021; Loyola-Gonzalez, 2019; Mahinpei et al., 2021). These black-box models embed hidden assumptions and biases (Suresh & Guttag, 2021), adding to the oversimplifications that arise during problem framing. The challenge is further amplified when AI models are embedded within larger, interconnected systems, where multiple models interact in ways that reduce transparency and make it harder to diagnose errors or unintended consequences. One response is to shift towards ‘white-box models’–such as decisions trees (Breiman et al., 1984) and rule-based systems (Loyola-Gonzalez, 2019), which offer greater interpretability. However, limiting RAD-AI research to white-box approaches risks overlooking valuable insights that could emerge from engaging critically with black-box methods and addressing their interpretability challenges in applied contexts.
@@ 2.3 Law, Policy, and Public Discourse
An increasing number of legal and policy developments are focusing on AI, with AI explicitly mentioned in laws passed by over 30 countries by 2024 (Maslej et al., 2024; UNESCO, 2024). Three main motivations are driving AI regulation: “addressing public problems”; “protecting, respecting or promoting human rights”; and creating the conditions to “achieve desirable futures” (UNESCO, 2024). Particular attention has been given to AI applications in areas that are considered “high risk”, including education, health, policing, and the workplace. Yet unlike traditional policymaking, tech policy tends to privilege a multistakeholder approach - directly involving the tech industry and its concerns in scoping regulation – with more marginal attention to social, economic, and environmental externalities (UNESCO, 2024).
@@
Engaging with law and policymaking processes and associated public debates provides researchers with opportunities to identify priorities for their own research and public engagement, often with great specificity (e.g., concern over facial recognition, data extraction, and high risk AI applications), rather than lofty principles and goals (Schiff et al., 2021). This should focus not only on AI-specific laws and policies, but particularly on established (and often extensive) legal requirements in the given application domain.
@@
Understanding national priorities and strategies around AI and application domains will be beneficial for understanding policymaking agendas and for researchers who rely on funding from national science agencies or foundations. RAD-AI provides a competitive edge in this context, as funding bodies increasingly stipulate both transdisciplinary and engaged research as component of fundings (Bednarek et al., 2025). The push to safeguard individual and community agency, along with building domestic digital capability and sovereignty, is already shaping regulation and funding globally (Chahal et al., 2022; Mügge, 2024; NSF, 2023), and must not be overlooked by the AI community.
@@
Maintaining public approval (often termed “social license”) is essential to the effective conduct of research. Communities generally expect a very high safety bar for new technologies, presenting a challenge for AI applications that require broad real-world testing to achieve market readiness (Hemesath & Tepe, 2023). For example, public apprehension has hindered the testing and deployment of autonomous vehicles (Hemesath & Tepe, 2023; Stilgoe & Cohen, 2021). Yet in application domains from food to pharamaceuticals to gene technology, both pre- and post-market testing are non-negotiable - there is no reason why applied AI should be any different.
@@
Strengthening the “connection between [AI’s] capability improvements and AI’s social or economic impacts” is critical (Narayanan & Kapoor, 2024). While methods-centric studies will prioritise a “narrow set of evaluation metrics” such as test loss and accuracy (Rolnick et al., 2024), they may overlook an innovation’s feasibility in solving real-world problems or its alignment with the societal values and priorities of the communities it serves. The “bottleneck for impact” often lies in the pace of product development and the rate of adoption, rather than methods advancement (Narayanan & Kapoor, 2024). Addressing these challenges requires deliberate efforts to align technical progress with societal needs, fostering both innovation and public trust.
2.4 Methods-Driven Research Cannot Account for Sectoral Nuances
@@
Nuances in Ethics. Applying AI to new fields of science carries considerable excitement. The 2024 Nobel Prize awards highlighted this transformative potential, showcasing AI advancements in the disciplines of Physics and Chemistry. However, as AI’s legitimacy and impact expand beyond computer science, the field is also interacting with disciplines that have strong histories of ethical standards, such as the Hippocratic Oath for physicians and the Obligation of the Engineer for engineers. These ethical standards guide the everyday workings of these professions and their fundamental societal obligations.
@@
It is important to recognise that effective engagement with nuances in ethics at these different levels will be messy. The integration of different fields has the potential to lead to deprioritisation or event displacement of existing priorities and ethics concerns. This has been seen in the Water, Sanitation, and Hygiene (WASH) field, where alignment with global health and development paradigms has obscured intrinsic aspects of the sector (de Wit et al., 2024). For example, prevailing paradigms prioritise universalisation (“creating global, mobile models”), responsibilisation (“locating responsibility with communities and individuals”), technicalisation (“rendering WASH actionable as a technical, depoliticized, global project”), and metricisation (“making problems and solutions measurable”) while overlooking localised values and the complex reality of lived experience in poverty (de Wit et al., 2024). More recent work has identified that in global water-climate nexus settings, the WASH priorities more aligned with climate change priorities have sidelined historic WASH priorities of equity and universal access to WASH. Similarly, in agriculture, applying ADML without intentional design could exacerbate historical and ongoing ethical issues, such as the use of child labour and the dispossession of smallholder or indigenous agricultural land (Hartman et al., 2022; Li, 2007; Tzachor et al., 2022).
@@
In agriculture, AI is increasingly seen as a component of an integrated suite of sociotechnical interventions needed to address food security. For example, global food production data, fundamental to tracking and responding to food security issues, is riddled with social and technical challenges that result in data scarcity worldwide (Kebede et al., 2024). Domain experts have suggested addressing this challenge in a nuanced way that involves thoughfully integrating AI into sociotechinical interventions (Kebede et al., 2024). In this sector, simply applying AI will not result in the integrated solutions that Kebede et al. argue will address the problem (2024).
@@
These examples point to the opportunity RAD-AI opens to work across sectors and with different types of actors to address contextual needs, responsibilities, and requirements.

### v2

@@ 1 Introduction
The artificial intelligence (AI) landscape is evolving rapidly, requiring researchers to engage with many aspects of AI integration into society. Innovations span diverse fields, from specific examples in medicine - where AI is helping with early detection of atrial fibrillation and providing real-time health insights in clinical settings (Briganti & Le Moine, 2020), to agriculture - where it promises to better predict crop yield (Elbasi et al., 2022), optimise fertiliser use (Elbasi et al., 2022), and analyse long-term agricultural trends (Hartman et al., 2022). Across both the public and private sectors, the widespread uptake of AI elevates the need for applied AI research to be responsive to contextual needs and societal values (Nissenbaum, 2009), ethical and legal considerations, and public concerns. Further, AI researchers have responsibilities to inform and educate the public regarding actual capabilities and limitations of AI, as an instrumental component of building trust in human-AI collaboration (McGrath et al., 2024).
@@
In recent years, the concept of Responsible AI (RAI) has gained significant traction in industry, government, and academic settings (Adams et al., 2024; Maslej et al., 2024; UNESCO, 2024). The concept of RAI (Confalonieri et al., 2021; Lu et al., 2023) has become more pronounced, as the failures, flaws, and harms of AI have been made visible (Cressie, 2023) due to the embedding of AI into everyday technologies. Responsible AI draws on prior technical work undertaken in the areas of fairness, transparency, and explainability, as means of responding to AI’s deficiencies (Confalonieri et al., 2021). Yet, two of the major weaknesses of RAI are that it lacks stable definition and that it fails, at core, to specify who is responsible, for what, to whom (Powles, 2023). Instead, RAI casts responsibility diffusely. Adams et al. (2024, p. 9), for example, define RAI as “the design, development, deployment and governance of AI in a way that respects and protects all human rights and upholds the principles of AI ethics through every stage of the AI lifecycle and value chain. It requires all actors involved in the national AI ecosystem to take responsibility for the human, social and environmental impacts of their decisions”.
@@
RAI gestures to earlier frameworks designed to ensure science and innovation are responsive to societal needs, including work on ethical, legal, and social implications, value sensitive design and, most significantly, Responsible Innovation (RI) or Responsible Research and Innovation (RRI). The classic formulation of RI advocates for a deliberative approach to research, centred on four dimensions: anticipation; reflexivity; inclusion; and responsiveness (Owen et al., 2013; Stilgoe et al., 2020). This model has been extensively applied and refined since the early 2010s, and adds considerable sophistication to RAI discourse, particularly around integrating affected communities into the research process (McCrea et al., 2024).
@@
This paper advocates for people-centred, participatory approaches in RAD-AI research, fostering collaboration, co-design, and efficacy-driven innovation beyond the constraints of purely method-driven approaches. By embedding AI development within real-world contexts and diverse perspectives, we can ensure that these technologies are not only technically robust but also socially and ethically viable. The question is: how do we get there?
@@ 2 From AI to RAD-AI: Framing AI Research as a Sociotechnical Challenge
2.1 Paradigms: From Technological Determinism to Centring People
The AI community needs to be wary of adopting the paradigm of technological determinism: the theory that technology inherently and autonomously changes society, unaffected by social pressures and constructs (Kline, 2015). Technological determinism can lead to AI researchers assuming that their work is inherently desirable, with the affordances of technology “becom[ing] ends in themselves, instead of the means for creating desired social ends” (Kline, 2015, p. 109). Instead of assuming, researchers should ask critical questions to ensure that their work does indeed align with contextual needs and societal values, shaping how technology functions, for whom, and to what effect (Davis, 2020). As Stilgoe and Cohen (2021) argue, by its nature, technological determinism precludes co-creation between those developing technologies and affected communities, since communities are the assumed eventual adopters of technology, rather than active players in how it is researched, developed, and shaped (Stilgoe & Cohen, 2021). The paradigm sees communities as “onlookers… displace[d] from the action” (Zenz & Powles, 2024, p. 3), incorrectly assuming that adoption is inevitable and public approval is not required (Stilgoe & Cohen, 2021; Zenz & Powles, 2024).
@@ 2.2 Breaking Out of AI (Black) Boxes
Problem Framing. One challenge we identify is the boundaries that experts impose when defining problems. Traditionally, professionals such as doctors, engineers, lawyers, and computer scientists, are taught to bound, map, characterise, and document a problem, situating it in a specific technical framing (Li, 2007). To distil problems into technical solutions, the socio-political complexities that define them are mostly glossed over, creating a skewed and incomplete approach to knowledge and in some cases even a “sidelining much of the data so painstakingly collected” (Li, 2007, p. 126). While expertise is important, it may narrow the frame of reference and leave out critical components of the complex world. What scientists initially hypothesise, collect data on, and build models around reflects existing power dynamics and interests, worldviews, and technical training, potentially missing critical nuances. We may miss understanding and nuance because it is not captured in the original framing of the question, the data collected, or the type of result being prioritised, including the specific metrics used to measure success (Rolnick et al., 2024). Addressing real-world problems with algorithms and AI systems requires looking beyond the constraints experts impose - both in how challenges are framed and in the broader contexts that shape them. These contexts, spanning history, politics, and power (Powles, 2023), not only define but also limit potential solutions.
@@
A second challenge is where the expert box is also bound by scientific reductionism – the scientist’s proclivity for oversimplification of reality (Greenhalgh, 2025). Concerns about oversimplification of AI approaches are evident across various fields, ranging from medicine (Sturmberg & Mercuri, 2024; Greenhalgh, 2025) to agriculture (Tzachor et al., 2022). For example, model emulation approaches (Gladish et al., 2018; Bolt et al., 2023; MacKinlay et al., 2025), such as those applied to complex agricultural models (Powell et al., 2023) offer crucial speedups and simplifications, effectively ‘jailbreaking’ their complexity. However, scepticism toward AI innovations often stems from the notion that AI seeks to replace, rather than complement and build off, well-established agricultural models. This kind of reductionism - which can involve excessive as well as insufficient parametrisation - generally overlooks the nuanced, interconnected, interdependent, and evolving nature of problems. It can also hide assumptions. As Sturmberg and Mercuri (2024, p. 3) note, “no problem exists in isolation” rather “all problems are embedded within a unique context.” Similarly, RAD-AI approaches will have nuanced differences depending on the context in which they are applied.
@@
Method Selection. A key challenge in AI method selection is the lack of interoperability of complex models. While ML approaches such as neural networks, support vector machines, random forests, and k-nearest neighbours can be tuned for high accuracy, they often obscure intermediate decisions, making it difficult to understand how conclusions are reached (Loyola-Gonzalez, 2019; Confalonieri et al., 2021; Mahinpei et al., 2021). These black-box models embed hidden assumptions and biases (Suresh & Guttag, 2021), adding to the oversimplifications that arise during problem framing. The challenge is further amplified when AI models are embedded within larger, interconnected systems, where multiple models interact in ways that reduce transparency and make it harder to diagnose errors or unintended consequences. One response is to shift towards ‘white-box models’– such as decisions trees (Breiman et al., 1984) and rule-based systems (Loyola-Gonzalez, 2019), which offer greater interpretability. However, limiting RAD-AI research to white-box approaches risks overlooking valuable insights that could emerge from engaging critically with black-box methods and addressing their interpretability challenges in applied contexts.
@@ 2.3 Law, Policy, and Public Discourse
An increasing number of legal and policy developments are focusing on AI, with AI explicitly mentioned in laws passed by over 30 countries by 2024 (Maslej et al., 2024; UNESCO, 2024). Three main motivations are driving AI regulation: to “address a public problem”; “protect, respect or promote fundamental and collective rights”; and create the conditions to “achieve a desirable future” (UNESCO, 2024, p. 41). Particular attention has been given to AI applications in areas that are considered “high risk”, including education, health, policing, and the workplace. Yet unlike traditional policymaking, tech policy tends to privilege a multi-stakeholder approach - directly involving the tech industry and its concerns in scoping regulation – with more marginal attention to social, economic, and environmental externalities (UNESCO, 2024).
@@
Engaging with law and policymaking processes and associated public debates provides researchers with opportunities to identify priorities for their own research and public engagement, often with great specificity (e.g., concern over face recognition, data extraction, and high risk AI applications), rather than lofty principles and goals (Schiff et al., 2021). This should focus not only on AI-specific laws and policies, but particularly on established (and often extensive) legal requirements in the given application domain.
@@
Understanding national priorities and strategies around AI and application domains will be beneficial for understanding policymaking agendas and for researchers who rely on funding from national science agencies or foundations. RAD-AI provides a competitive edge in this context, as funding bodies increasingly stipulate both transdisciplinary and engaged research as component of funding (Bednarek et al., 2025). The push to safeguard individual and community agency, along with building domestic digital capability and sovereignty, is already shaping regulation and funding globally (Chahal et al., 2022; NSF, 2023; Mügge, 2024), and must not be overlooked by the AI community.
@@
Maintaining public approval (often termed ’social license’) is essential to the effective conduct of research. Communities generally expect a very high safety bar for new technologies, presenting a challenge for AI applications that require broad real-world testing to achieve market readiness (Hemesath & Tepe, 2023). For example, public apprehension has hindered the testing and deployment of autonomous vehicles (Stilgoe & Cohen, 2021; Hemesath & Tepe, 2023). Yet in application domains from food to pharmaceuticals to gene technology, both pre- and post-market testing are non-negotiable - there is no reason why applied AI should be any different.
@@
Strengthening the “connection between [AI’s] capability improvements and AI’s social or economic impacts” is critical (Narayanan & Kapoor, 2024). While method-centric studies will prioritise a “narrow set of evaluation metrics” such as test loss and accuracy (Rolnick et al., 2024, p. 2), they may overlook an innovation’s feasibility in solving real-world problems or its alignment with the societal values and priorities of the communities it serves. The “bottleneck for impact” often lies in the pace of product development and the rate of adoption, rather than methods advancement (Narayanan & Kapoor, 2024). Addressing these challenges requires deliberate efforts to align technical progress with societal needs, fostering both innovation and public trust.
2.4 Method-Driven Research Cannot Account for Sectoral Nuances
@@
Nuances in Ethics. Applying AI to new fields of science carries considerable excitement. The 2024 Nobel Prize awards highlighted this transformative potential, showcasing AI advancements in the disciplines of Physics and Chemistry (Li & Gilbert, 2024). However, as AI’s legitimacy and impact expand beyond computer science, the field is also interacting with disciplines that have strong histories of ethical standards, such as the Hippocratic Oath for physicians and the Obligation of the Engineer for engineers. These ethical standards guide the everyday workings of these professions and their fundamental societal obligations.
@@
It is important to recognise that effective engagement with nuances in ethics at various levels will be messy. The integration of many fields has the potential to lead to deprioritisation or event displacement of existing priorities and ethics concerns. This has been seen in the Water, Sanitation, and Hygiene (WASH) field, where alignment with global health and development paradigms has obscured intrinsic aspects of the sector (de Wit et al., 2024). For example, prevailing paradigms prioritise universalisation (“creating global, mobile models”), responsibilisation (“locating responsibility with communities and individuals”), technicalisation (“rendering WASH actionable as a technical, depoliticised, global project”), and metricisation (“making problems and solutions measurable”) while overlooking localised values and the complex reality of lived experience in poverty (de Wit et al., 2024, p. 5). Recent work indicates that in global water-climate contexts, the growing emphasis on climate-aligned WASH priorities has sidelined historic WASH priorities of equity and universal access to WASH (Cullen et al., 2023). Similarly, in agriculture, applying ADML without intentional design could exacerbate historical and ongoing ethical issues, such as the use of child labour and the dispossession of smallholder or indigenous agricultural land (Li, 2007; Hartman et al., 2022; Tzachor et al., 2022).
@@
In agriculture, AI is increasingly seen as a component of an integrated suite of sociotechnical interventions needed to address food security. For example, global food production data, fundamental to tracking and responding to food security issues, is riddled with social and technical challenges that result in data scarcity worldwide (Kebede et al., 2024). Domain experts have suggested addressing this challenge in a nuanced way that involves thoughtfully integrating AI into sociotechnical interventions (Kebede et al., 2024). In this sector, simply applying AI will not result in the integrated solutions that Kebede et al. (2024) argue will address the problem.
@@
These examples point to the opportunity RAD-AI opens to work across sectors and with many types of actors to address contextual needs, responsibilities, and requirements.